OpenStudy (hockeychick23):

Floyd is playing a game where he is trying to roll a 3 with a standard die. If he gets a 3 in any of his first 8 rolls, he wins; otherwise, he loses. What is the probability that Floyd loses the game?

OpenStudy (tkhunny):

What is the probability of a 3 an any single roll? What is the probability of getting NO 3 in 8 rolls?

OpenStudy (hockeychick23):

1/6 in 1 roll and none maybe 1 in 48?

OpenStudy (anonymous):

There are 6 posibble options when you roll a die : 1,2,3,4,5 and 6 yes 1/6 that when you roll you will get a 3 5/6 that when you roll you wont get a 3

OpenStudy (hockeychick23):

yepp, so then what do i do with the 8 now that i have the 1/6 and 5/6?

OpenStudy (tkhunny):

No guessing. p(3) = 1/6 p(not 3) = 1-p(3) = 1-(1/6) = 5/6 p(no 3 in 8 rolls) = ????

OpenStudy (tkhunny):

Absolutely not. 0 <= p(stuff) <= 1

OpenStudy (tkhunny):

Independent events. Use the multiplication rule.

OpenStudy (hockeychick23):

ok so 1/6*1/6*1/6

OpenStudy (anonymous):

dang it that was incorrect .... \[p^n\] where p is the probability that you wont roll a 3 and n is the amount of times you can roll the die

OpenStudy (hockeychick23):

got it, 23.3, thanks!

OpenStudy (anonymous):

.233 watch your decimals

OpenStudy (tkhunny):

Not quite. Heartbreak had the compliment of the probability of rolling a 3 all eight time. How may times are we rolling the die? You seem to have just 3. And we don't care about rolling 8 3s. Just one. Thus, we need to concern ourselves with rolling NO 3s. p(3) = 1/6 p(not a 3) = 5/6 p(no 3 in 8 tries) = (5/6)^8 Come on! You're making me do all the work.

OpenStudy (hockeychick23):

sorry i meant 23.3%

OpenStudy (anonymous):

Yup thats correct

OpenStudy (hockeychick23):

thankss(: